OpenStudy (anonymous):

Which number has a factor of 3? Choose one answer. a. fifty-two b. seventy-six c. one hundred twenty-five d. one hundred thirty-five

OpenStudy (beginnercountrysinger):

its D

OpenStudy (whpalmer4):

@BeginnerCountrySinger please explain, don't just give answers.

OpenStudy (whpalmer4):

52=2*26=2*2*13 76=2*38=2*2*19 125=5*25=5*5*5 135=3*45=3*3*15=3*3*3*5

OpenStudy (beginnercountrysinger):

Sorry! when you break down the factors of 135 you get 3 and 45. its easiest to divide the number by the factor c:

OpenStudy (anonymous):

thank you :D I am not very good at the whole factors and primes and stuff ^U^

OpenStudy (beginnercountrysinger):

if you need help tag me and i'll explain if i can c:

OpenStudy (whpalmer4):

some easy tests for divisibility: any even number (ends with 0,2,4,6,8) is divisible by 2 any number divisible by 3 has a sum of digits which add up to a number divisible by 3. For example, 135 has digits 1,3,5 and 1+3+5 = 9, which is divisible by 3, so 135 is divisible by 3. If you get a multi-digit sum of digits, just repeat the process: 9999 has digits 9,9,9,9 and digit sum 9+9+9+9 = 36. 36 has digits 3,6, and 3+6 =9 which is divisible by 3, so 9999 is divisible by 3. any number ending with 0 or 5 is divisible by 5 The first handful or two of prime numbers: 2,3,5,7,11,13,17,19,23,29,31